OpenStudy (anonymous):

Help, medal will be awarded! Point A is located at (−2, 2), and D is located at (4, −2). Find the coordinates of the point that lies halfway between A and D. (−1, 1) (0, 1) (1, 0) (1, 1)

OpenStudy (haseeb96):

just use the mid point formula here (x,y) = (x2+X1/2 , y2 +y1 /2 ) here (x1,y1)=(-2,2) (x2,y2) =(4,-2)

OpenStudy (haseeb96):

C is the correct answer now try to do this question yurself okay @izzyvelazco

OpenStudy (anonymous):

i am not taking this class sorry haha, I'm trying to finish all these assignments today for my friend

OpenStudy (bradely):

Calculate the midpoint ((-2+4)/2,(2+(-2))/2)) (2/2,0,2) =(1,0) option 3 correct Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

thank you both though you're a huge help!

OpenStudy (haseeb96):

i am going to pray so i will come after 15 mins okay

OpenStudy (haseeb96):

just wait then i will help u in solving all the questions

OpenStudy (anonymous):

okay!